Two light waves of the same wavelength travel in the same direction in the same region of space. One has an intensity of 5I0 and the other has an intensity of 13I0. What is the total intensity of the light in this region if...
(a) the waves are incoherent
(b) the waves are coherent and in phase (c) the waves are coherent and 180° out of phase
Answer #1

let I1 = 5*Io
I2 = 13*Io

a) when the waves are incoherent,

the total intensity lies between I2 + I1 and I2 - I1

I2 + I1 = 13*Io + 5*Io = 18*Io

I2 - I1 = 13*Io - 5*Io = 8*Io

so, total intensity lies between 8*Io and 18*Io

b) when the waves are coherent and in phase

I_total = I1 + I2

= 5*Io + 13*Io

= 18*Io

c) when the waves are coherent and 180° out of phase

I_total = I2 - I1

= 13*Io - 5*Io

= 8*Io